From GlueXWiki
Jump to: navigation, search

May 3, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • Production status
    • Third package refurbishment status
    • Third package tests (Lubomir)
  2. Engineering update (Bill)
  3. Electronics update (Chris)
  4. Tests with first package at 126 FDC E-log (Beni)
  5. Other


Participants: Bill, Dave, Chris, Nick, Simon, Beni, and Lubomir.


- Dave: for package #3: wire planes 4 & 13 were deadened, now all the elements for this package are ready. Package #4: wire planes 24 and 25 ready for deadening, will be done on Monday. Chris will put components on wire plane 26 tomorrow. Anatoly is gluing daughter cards on cathode 47. Tina working on cathodes 49 and 50. Casey is filling the groove on the second type-1 cathode to be converted into type-2.

- Now we have four cells in package #3 installed. After the first cell installed and tested we put the next three cells on Wednesday, Thursday and Friday at once without tightening the chamber in between. On Friday after starting the gas we found huge leakage from the top spacer ring. It turned out the O-ring was smashed at one place. We suspect it happened because of the grease and incorrect use of the installation metal pins. When you insert the pins up into the new holes during the next cell installation, the pin can lift some of the frames and create gaps in between them. When you have gap the O-ring can stick to the surface above the groove (due to the grease), and then get out of the groove. To exclude this from happening again, when inserting the metal pins we should always apply force from the top (by fingers) to avoid separation of the frames. It is dangerous because you can damage an O-ring anywhere in the already installed cells which will require re-installation of the whole package. In case we have more problems Bill will consider designing a tool for pulling the pins. The damaged O-ring was replaced on Friday and we started flushing the package with gas.

- With four cells in the third package after 6 days flushing with gas we have 130ppm oxygen with 220 sccpm gas flow. The gas leakage that we had before was significantly reduced by putting a 5mil shim in between the bottom gusset ring and the end window. Now the leakage detector shows between upper 10^-4 and lower 10^-5. HV was applied on the top cell without problems (2100V) and all the channels were tested with a source. Today we continue with the tests of the next cell.


- Installation of the remaining two cells of the third package will require new metal pins. The pins that are now there are bent and can't go up. Bill ordered them and we will have them at the beginning of the new week. We discussed if we need better (and more expensive) pins: as long as we have spares we will be fine with these.


- Chris is looking for the parts needed for the grounding. Beni proposed a way for local connection between the grounds of the wire frames and the two neighboring cathodes: using copper tape (not adhesive), Rohacell piece inside the wire frame/cathode gap at the periphery, and a 1cm clamp (~3/4inch wide) that holds all together. Chris got some samples, and asked for the clamp that we need. In addition there will be inter-cell connections at the six places between the fast-ons on the wire frames. As Beni found using foil gives better grounding than with wires, not clear why. In about three weeks we will need the grounding material, Beni will show us the places to install this ground on the third package.

- Chris checked the dimensions of the first article LV cables, everything is fine except for the labeling that will be corrected. After the meeting Beni and Lubomir connected the LV cables at several critical places in the first package in 126: the sizes are OK, but it was very difficult to do the connections even using Bill's tools, mainly because of the tubing. Nick is working on the LV distribution boards, will attached on the LV rack and will have one per package for the cathodes and one for all the anodes.

Tests with first package at 126

- Beni did his local grounding at hundred(?) places in the first package and reduced the threshold from 13 to 4.5V! He showed scope pictures (see FDC elog ID 59 and 60 linked above): the noise is below 2mV! Strong after-pulsing visible in the cathode signals: with this 90/10 gas probably the HV is too high. Beni will check the same on the wires but needs analog preamp card.

- With the lower threshold the efficiencies on the chamber are higher (80-100%) (see Beni's elog). First estimation of the anode resolution is ~300 microns averaged over all the distances to the wire, worse than it was in the prototype. Now Beni uses the top and the third from the top cells, since they have parallel wires. Bill corrected Beni about the distance between the cells used in this calculations, hopefully this will improve the resolution.